OpenStudy (anonymous):

find the number of ways 3 parties can be assigned to 3 of the 12 empty booths of a restaurant

OpenStudy (anonymous):

@mathstudent55

jimthompson5910 (jim_thompson5910):

Let A, B, C be the 3 parties and D be considered an empty party or table so one combo is ABCDDDDDDDDD and another combo is ACDDDDDBDDDD etc etc

OpenStudy (anonymous):

soo the answer is ?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

there are 12 P 12 = 12! = 479001600 ways to arrange the 12 positions....BUT...D repeats itself (there are 9 copies of D) so you have to divide by 9! to get 479001600/(9!) = 479001600/362880 = 1320

jimthompson5910 (jim_thompson5910):

so there are 1320 different ways to assign parties A, B, C to the 12 tables

OpenStudy (anonymous):

thanks!